OpenStudy (kearsten):

solve by elimination, x+y=7 and 3x-2y=11. I don't know whether I'm supposed to add or subtract. show steps plz.

OpenStudy (kearsten):

If i knew whether i add or subtract i could get it done my self but i get confused with the addition and subtraction part.

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} x+y=7&{\color{brown}{ 2}}\implies &2x\cancel{+2y}=14\\ 3x-2y=11&\implies &3x\cancel{-2y}=11 \\\hline\\ &&\square ?+\square ?=\square ? \end{array} }\)

OpenStudy (wolf1728):

A) x+y=7 B) 3x-2y=11 You have to look to see which unknown an be eliminated by addition or subtraction. Neither one can be eliminated so do you know what to do next?

OpenStudy (kearsten):

not really

OpenStudy (jdoe0001):

you could, think of it as just addition if you add a "negative number", then you're still adding, kinda 3 + 3 = 6 3 - 2 = 1 or 3 + (-2) = 1, same thing

OpenStudy (kearsten):

where do u get the other 3? it 1x and 3x

OpenStudy (jdoe0001):

hmmm that was just an example meaning that, a subtraction could also be seen as an addition, with a negative number

OpenStudy (kearsten):

ur confusing me

OpenStudy (wolf1728):

Hello kearsten - guess jdoe0001 left I can take over

OpenStudy (jdoe0001):

hmm well... notice above... the elimination, we used a factor of "2" so, "add" or "subtract" away then

OpenStudy (kearsten):

ok

OpenStudy (wolf1728):

As I previously posted: A) x+y=7 B) 3x-2y=11 You have to look to see which unknown an be eliminated by addition or subtraction. Neither one can be eliminated so do you know what to do next?

OpenStudy (kearsten):

idk, my question says to use elimination.

OpenStudy (wolf1728):

You have to eliminate the x or the y by adding or subtracting the equations. You can see it can't be done just by addition or subtraction right?

OpenStudy (kearsten):

ok

OpenStudy (wolf1728):

Let's look at the equations again A) x+y=7 B) 3x-2y=11 We will have to multiply one of those by a number so that when the 2 are added or subtracted, we can eliminate the x or y

OpenStudy (kearsten):

so times equation a by 3 or 2?

OpenStudy (wolf1728):

If you multiply equation A by 2 you'll see it produces 2y equation B has a -2y so adding those equations will eliminate the y

OpenStudy (kearsten):

so equation a will be 2x+2y=7? or would it be x+2y=7?

OpenStudy (wolf1728):

you have to multiply ALL of the equation by 2

OpenStudy (wolf1728):

neither of those equations you typed are correct

OpenStudy (kearsten):

so 2x+2y=14?

OpenStudy (wolf1728):

yes !!

OpenStudy (wolf1728):

Now add it to equation B

OpenStudy (kearsten):

x=5?

OpenStudy (wolf1728):

Let's not skip anything 2x + 2y = 14 3x-2y=11 5x = 25 x=5

OpenStudy (kearsten):

ok i got that but simplified

OpenStudy (wolf1728):

the next step is to get the value of x=5 and put it into either equation

OpenStudy (kearsten):

3x-2y=11 3(5)-2y=11 15-2y=11 2y=26 y=13?

OpenStudy (kearsten):

or is it x+y=7 5+y=7 y=2?

OpenStudy (wolf1728):

Not quite 15-2y=11 4 = 2y Y=2

OpenStudy (wolf1728):

Yes y = 2

OpenStudy (kearsten):

awesome, i got it right! sorta. thanks for the help wolf!

OpenStudy (wolf1728):

glad to help :-)

OpenStudy (kearsten):

??? no idea y that change spots. thnks for the help ur a life saver